find fa,b

Fa, b= K garb Chat Parb 13 Pa 16 = Cxa-x) 2+ Cya-4b) Ý yb=0 Xq = 12.5 e 2m ya-21.76 am Xp= 25 e 2m Xc=0
Answer #1

- oc 12.5x10-2 î + 211771629 25x162 i to Za = 12.5x152 - 25x162 9 +2 1.79 - of = (-1205x1622 +21-7 9 11x1027 = (-1225 +21.79)